# When do Neural Networks Learn World Models?

### v2

@@ C.1 Technical Lemmas
Proof of (i). If $U$ is compatible, then by Lemma C.14 we have $\mathrm{deg}_{U}(f)=\mathrm{deg}(f)$ for any Boolean function $f$. This immediately gives the equivalence between (71) and (56) and hence their minimizers. We thus have that $\Phi^{*}$ learns the world model up to negations and permutations as in Theorem 4.8.
@@
Appendix D Experiment Details
This section presents additional experiment details. All of our experiments were conducted using PyTorch (Paszke et al., 2019) and on NVIDIA V100, NVIDIA A100, and NVIDIA H100 GPUs.
D.1 Polynomial Extrapolation
@@
D.2 Learning Physical Laws
@@
We train a decoder-only transformer (Vaswani et al., 2017) conditioned on the first 3 frames to predict the remaining frames (expect for the timestamp dimension). For our method, we simply replace every MLP in the original transformer with our modified MLP as in Section D.1 (replacing ReLU by GELU). We use teacher-forcing in training that is similar to next-token prediction, i.e., the model only needs to predict the next frame (starting from the 4-th frame to the last 32-th frame) given all ground-truth frames prior to it. We use the following model hyperparameters:
@@
Appendix E Additional Results and Discussion
@@ E.1 Polynomial Extrapolation
E.2 Learning Physical Laws

### v3

@@ E.1 Technical Lemmas
Proof of (i). If $U$ is compatible, then by Lemma E.14 we have $\mathrm{deg}_{U}(f)=\mathrm{deg}(f)$ for any Boolean function $f$. This immediately gives the equivalence between (71) and (56) and hence their minimizers. We thus have that $\Phi^{*}$ learns the world model up to negations and permutations as in Theorem 4.8.
@@
Appendix F Experimental Details
This section presents additional experimental details. All of our experiments were conducted using PyTorch (Paszke et al., 2019) on NVIDIA V100/A100 GPUs.
F.1 Polynomial Extrapolation
@@
F.2 Learning Physical Laws
@@
We train a decoder-only transformer (Vaswani et al., 2017) conditioned on the first 3 frames to predict the remaining frames (expect for the timestamp dimension). For our method, we simply replace every MLP in the original transformer with our modified MLP as in Section F.1 (replacing ReLU by GELU). We use teacher-forcing in training that is similar to next-token prediction, i.e., the model only needs to predict the next frame (starting from the 4-th frame to the last 32-th frame) given all ground-truth frames prior to it. We use the following model hyperparameters:
@@
Appendix G Additional Results and Discussion
@@ G.1 Polynomial Extrapolation
G.2 Learning Physical Laws